The soles of a popular make of running shoe have a force constant of 2.3×105 N/m. Treat the soles as ideal springs for the following questions.

Part A: If a 62 kg person stands in a pair of these shoes, with her weight distributed equally on both feet, how much does she compress the soles?

X = ________ m

Part B: How much energy is stored in the soles of her shoes when she's standing?

U = ________ J
Answer #1

Solution)

Given,

Part a)

Force constant = 2.3×10^5 N/m

compression on soles = (0.5 * 62 * 9.8)/( 2.3×10^5)

= 0.00132 m

========

Part b)

Energy stored in 1 shoe = 1/2 * 2.3×10^5 * (0.00132)^2

= 0.2003 J

Hence, Energy stored in both shoes = 2* 0.2003 J

= 0.4007 J

===========